PostScript

Information Sources

1 Books

• Blue Book: PostScript Language Tutorial and Cookbook, Adobe Systems Incorporated, 1985.

• Red Book: PostScript Language Reference Manual, Adobe systems Incorporated, 1985.

2 /mit/postscript

/mit/postscript is a world accessible file system for sharing PostScript hacks. There are lots of various PostScript sources there for people to copy and use. Additionally, the file system is world writeable so you can add your own PostScript hacks. To access, simply attach postscript.

N.B. /mit/postscript/quick-ref.PS is an alternate quick reference for PostScript.

3 pshacks

Pshacks is a discuss meeting for discussion about PostScript. If you’ve never used discuss, stop by the SIPB office (11-205) and ask for a copy of Using Discuss. To add meeting, issue the following command from discuss:

am bloom-beacon.mit.edu:/usr/spool/discuss/pshack

4 comp.lang.postscript

Usenet netnews carries some PostScript relevant discussions in comp.lang.postscript. To access this, simply add comp.lang.postscript to your subscription list by issuing the command g comp.lang.postscript. If you’ve never used netnews before stop by the SIPB office (11-205) and ask for a copy of Netnews: A One-Page Guide to Usenet.